Suppose that the average number of cars abandoned weekly on a certain highway is2.2.


Approximate the probability that there will be:


(a) no abandoned cars in the next week;


(b) at least 2 abandoned cars in the next week. . . " .....


(c) State the key assumptions needed to' make your answers in a) and b) reasonable

1
Expert's answer
2022-05-23T22:54:13-0400

a)

"X\sim Po(\lambda)"

"\lambda=2.2"


"P(X=0)=\dfrac{e^{-2.2}(2.2)^0}{0!}=0.110803"

b)

"X\sim Po(\lambda)"

"\lambda=2.2"


"P(X\ge 2)=1-P(X=0)-P(X=1)"

"=1-\dfrac{e^{-2.2}(2.2)^0}{0!}-\dfrac{e^{-2.2}(2.2)^1}{1!}"

"=0.645430"



(c) if "n" is large and "p" is close to "0," the Poisson distribution can be used,

with "\u03bc = np," to approximate binomial probabilities.